The perimeter of the state of Wyoming is 1280 mi. The width is 90 mi less than the length. Find the width and the length.
Answer by mananth(16946)

let length be x mi
width = x-90 mi
..
perimeter = 2*(L+W)
1280 = 2*(x+x-90)
1280=2*(2x-90)
1280/2 = 2x-90
640 = 2x-90
640+90=2x
730=2x
x= 365 miles the length
..
width = 365 -90 = 275 miles
CHECK
275*2+365*2=550+730=1280